Rick has zombies in his barn. On Monday he started adding 2 more zombies every day. On day 10 there were 30 zombies and his barn. Write an equation to model. how many zombies were in the barn to start?
Answer by Boreal(15235)

an=a1+(n-1)d
an is day 10 so a10
a10=a1+(10-1)2=30
12=a1 , the number on day 1